OpenStudy (anonymous):

FREE Metal!! Evaluate the expression described below if the variable given is 9. Two times the difference of one-third of a number and three. A. –9 B. –3 C. 0 D. 12

OpenStudy (anonymous):

C? 1/3(9)=3 3-3=0 0*2=0 I could be wrong but it makes sense...